OpenStudy (18jonea):

Graph y = 4x – 3 as a translation of its parent function.

OpenStudy (18jonea):

@Mehek14

OpenStudy (18jonea):

@sleepyjess

OpenStudy (18jonea):

@ILovePuppiesLol

ILovePuppiesLol (ilovepuppieslol):

parent function?

OpenStudy (18jonea):

yes

OpenStudy (18jonea):

@ILovePuppiesLol

OpenStudy (18jonea):

@Preetha

OpenStudy (18jonea):

@myininaya

OpenStudy (18jonea):

@Michele_Laino

myininaya (myininaya):

Is there a stretch from the parent function? Is there a shift from the parent function? Is there a reflection from the parent function to this function?

OpenStudy (18jonea):

How would i know that?

myininaya (myininaya):

\[y=f(x) \text{ let's call this the parent function } \\ \] \[\text{ Let } c>0 \\ y=f(x+c) \text{ shifts the graph of } f(x) \text{to the left } c \text{ units } \\ y=f(x-c) \text{ shifts the graph of } f(x) \text{ to the right } c \text{ units }\] \[y=f(x)+c \text{ shifts the graph of } f(x) \text{ up } c \text{ units } \\ y=f(x)-c \text{ shifts the graph of } f(x) \text{ } c \text{ units down }\] so do you think you have any kind of shift? also do you know the parent function?

OpenStudy (18jonea):

down

myininaya (myininaya):

\[y=K f(x) \text{ then } \\ \text{ if } K>1 \text{ then the graph of } f(x) \text{ is stretched in the } y-\text{ direction } \\ \text{ if } 0<K<1 \text{ the graph of } f(x) \text{ is compressed }\]

myininaya (myininaya):

If K is negative that will involved a reflection too

myininaya (myininaya):

So do you know the parent function of y=4x-3 ?

myininaya (myininaya):

hmm... none of those are the answer...

OpenStudy (18jonea):

no i dont im still kinda confused

OpenStudy (18jonea):

Thats why im confused

myininaya (myininaya):

just so you know you posted "Graph y=4x-3 ..."

myininaya (myininaya):

this is a line none of those are lines

OpenStudy (18jonea):

Graph y = 4^x – 3 as a translation of its parent function. sorry the x is an exponent

OpenStudy (18jonea):

@myininaya

myininaya (myininaya):

do you know what the graph of y=4^x looks like?

myininaya (myininaya):

if not you might remember what the graph of y=e^x looks like?

myininaya (myininaya):

draw either and move the graph 3 units down as you suggested earlier what graph in your options is most close to what you drew ?

OpenStudy (18jonea):

so what option would it be

myininaya (myininaya):

yep that works

OpenStudy (18jonea):

so what option would it be? for the parent

myininaya (myininaya):

or we suppose to graph the parent function or graph y=4^x-3 using the parent function?

myininaya (myininaya):

the directions say graph y=4^x-3 ....

OpenStudy (18jonea):

Graph y = 4^x – 3 as a translation of its parent function.

myininaya (myininaya):

the parent function looks like this: 

myininaya (myininaya):

if you drag that down 3 units what do you get

myininaya (myininaya):

the graph you posted

OpenStudy (18jonea):

ok thank you

OpenStudy (18jonea):

can you help me with a few more problems?

myininaya (myininaya):

maybe

OpenStudy (18jonea):

ok want me to just keep this post going?

OpenStudy (18jonea):

@myininaya

myininaya (myininaya):

sorry someone else hit me up just now

OpenStudy (18jonea):

ok so you cant help or you can?

OpenStudy (18jonea):

can you just help me really quick

myininaya (myininaya):

we can skip a lot of work if you know that f(g(x))=x where f and g are inverses

OpenStudy (18jonea):

ok so what would my answer be?

myininaya (myininaya):

ok if you don't know how to use that then find the inverse function of g

myininaya (myininaya):

that is solve \[y=\sqrt[3]{x+1}\] for x

OpenStudy (18jonea):

what is the inverse?

OpenStudy (18jonea):

I dont really know how to solve that

OpenStudy (18jonea):

@myininaya

myininaya (myininaya):

can you try cubing both sides and see what happens